An object is moving along x-axis in such a way that its displacement is given by x=30+20t2 ,where x in meters and t in seconds .what will be initial and final velocities.
1
Expert's answer
2012-03-30T07:06:35-0400
Let's find the velocity of an object:

V = x' = (30+20t²) = 2*20t = 40t.

Let T0 be the initial time and T1 be the final time. Then initial and final velocities will be

V0 = 40·T0

and

V1 = 40·T1,

respectively.